To compare customer satisfaction levels of two competing ice cream companies, 8 customers of Company 1 and 5 customers of Company 2 ware randomly selected and were asked to rate their ice creams on a five-point scale, with 1 being least satisfied and 5 most satisfied. The survey results are summarized in the following table: Company 1 Company 2 ๐‘›1=8 ๐‘›2=5 ๐‘‹1 ฬ…ฬ…ฬ…ฬ…=3.21 ๐‘‹2 ฬ…ฬ…ฬ…ฬ…=2.22 ๐‘†1= 0.31 ๐‘†2= 0. 21 By assuming both equal variances and unequal variances: a. Construct an interval estimate at 95ยฐ/ confidence interval for difference of means of two ice cream companies.


1
Expert's answer
2021-01-13T16:57:15-0500

We have that

n1 = 8

"\bar x_1 = 3.21"

s1 = 0.31

n2 = 5

"\bar x_2 = 2.22"

s2 = 0.21


Assuming both variances are equal:

Since samples are less than 30 in this problem we are dealing withย t-distirbution with n1ย + n2ย โ€“ 2 = 8 + 5 โ€“ 2 = 11 degrees of freedom.

The table t-value for a 95% confidence interval with 25 df is t0.025, 11 = 2.201

The formula forย a 95% confidence interval for the difference of two population means:


"(\bar x_1-\bar x_2)\pm t_{0.025, 11}s_p\sqrt{\frac{1}{n_1}+\frac{1}{n_2}}"

where

"s_p=\sqrt{\frac{s_1^2(n_1-1)+s_2^2(n_2-1)}{n_1+n_2-2}}"

"s_p=\sqrt{\frac{0.31^2\cdot7+0.21^2\cdot4}{8+5-2}}=0.28"

Confidence interval:

"(3.21-2.22)\pm2.201\cdot0.28\sqrt{\frac{1}{8}+\frac{1}{5}}=0.99\pm0.35"

We are 95% confident that the difference in the two population means is between 0.64 and 1.34. Zero is not in this interval so there is a significant difference of means of two ice cream companies.


Constructing a confidence interval for the difference of means when both variances are not equal:


"(\bar x_1-\bar x_2)\pm t\sqrt{\frac{s_1^2}{n_1}+\frac{s_2^2}{n_2}}"

where df is the smaller of n1โ€“1 and n2โ€“1

df = min(8-1, 5-1)=4

Therefore the table t-value for a 95% confidence interval with 4 df is t0.025, 4 = 2.776

"(3.21-2.22)\pm 2.776\sqrt{\frac{0.31^2}{8}+\frac{0.21^2}{5}}=0.99\pm 0.4"

We are 95% confident that the difference in the two population means is between 0.59 and 1.39. Zero is not in this interval so there is a significant difference of means of two ice cream companies.